Jacque has 3 L of a 50% antifreeze mixture. How much is 75% mixture should be added to get a mixture that is 60% antifreeze?
Answer by richwmiller(17219)

0.75*x+0.5*3=0.6(3+x)
0.75*x+1.5=1.8+0.6x
0.75x-0.6x=1.8-1.5
0.15x=0.3
x=2 liters at 75% is added to 3 liters of 50% to get 5 liters at 60%
check
0.75*2+0.5*3=0.6(3+2)
1.5+1.5=0.6(5)
3=3
ok